Let’s do a numerical example of how to find a good’s elasticity. A grocery store manager has raised the price of yellow peppers from $3 to $4. As a result, he notices the sales of yellow peppers have fallen from 500 to 300. What is the elasticity value of yellow peppers using the formula for price elasticity of demand?
Answer #1

P0=3 And P1=4

Q0=500 and Q1=300

Elasticity of Demand=(change in Quantity Demand/change in price)×(Price/Demand)

Ed=(200/1)×(3/500=600/500=-1.2